OpenStudy (anonymous):

At the end of The Wizard of Oz, the Scarecrow earns his diploma. He then says, “The sum of the square roots of any two sides of an isosceles triangle is equal to the square root of the remaining side”. How should this statement of the Pythagorean Theorem correctly read?

OpenStudy (anonymous):

A. The sum of the squares of any two sides of a right triangle is equal to the square of the shortest side.. B. The sum of the square roots of any two sides of an isosceles triangle is equal to the square root of the largest side. C. The difference of the squares of any two legs of a right triangle is equal to the square of the hypotenuse. D. The sum of the squares of any two legs of a right triangle is equal to the square of the hypotenuse.

OpenStudy (anonymous):

The Wiz incorrectly stated the theorem like:\[\sqrt{a} + \sqrt{b} = \sqrt{c}\]

OpenStudy (anonymous):

so would my answer be C ?

OpenStudy (anonymous):

I don't think so... Pythagorean theorem is usually stated as a sum, not a difference.

OpenStudy (anonymous):

i'd have go to with B

OpenStudy (anonymous):

am i wrong again ? /:

OpenStudy (anonymous):

you only have 4 choices ;) Are you just working through them all? B talks about isosceles triangles... but Pythagorean talks about right triangles.

OpenStudy (anonymous):

Have you seen Pythagorean theorem stated like: \[a ^{2} + b ^{2} = c^{2}\]

OpenStudy (anonymous):

yes i have

OpenStudy (anonymous):

reading the answers more and more , D sounds like it fits the best out of them all to me . it seems to make more sense .

OpenStudy (anonymous):

Let's look at D.... I think it describes that, but you have to be careful... the wording in these is a little tricky...

OpenStudy (anonymous):

Sum of squares... that is the "a^2 + b^2" part Right triangle... got it... = square of hypotenuse... that's the c^2 part

OpenStudy (anonymous):

D looks like it works....

OpenStudy (anonymous):

thank you for helping me out ! much appreciated !

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

(: